# BIOPTIC - a target-agnostic potency-based small molecules search engine

### v2

@@ 4.1 Modeling
Our model is a SMILES-based transformer [21], which we train in two phases. The Byte-Pair Encoding (BPE) tokenization [22] is applied to SMILES strings at each phase. During the first phase, the model is pre-trained in an unsupervised fashion on a large corpus of unlabeled molecular data from PubChem and Enamine REAL Space. The training is performed via a masked language modeling following the procedure from RoBERTa [23] (see fig. 1a). This phase allows the model to learn the SMILES grammar.
In the second stage, the transformer model is augmented with a multi-pooling module that aggregates hidden layer output in three different ways: taking classification token output as is, and taking maximum and average along the sequence axis of all the remaining tokens. These aggregated vectors are passed into their corresponding linear layers for projection to a lower dimension following up with concatenation to form the final embedding. This embedding is further normalized and passed to the classification layer where the task is to predict which targets a given molecule is active to. Here, the model is trained on BindingDB. As the majority of labels are not available for the BindingDB drug-target pairs, we mask them out to not introduce additional noise into the training (fig. 1b).
@@ 4.2 Optimization
The chosen model is RoBERTa, with a vocabulary size of 500, 6 hidden layers with a hidden size of 384, 8 attention heads, and an intermediate size of 1024, while the rest of the parameters are as described in [23]. The model has a total of 8.7 million parameters. Initially, it is pre-trained in an unsupervised fashion on a large corpus of data comprising 115 million unique molecules from PubChem and a random set of 48 million molecules from the Enamine REAL Space library, totaling over 160 million molecules. The training is performed using masked language modeling following the procedure from RoBERTa [23] (fig. 1).
@@ 4.3 Ultra-fast molecules retrieval system
For the Enamine REAL Space of 40 billion molecules, the inference of the neural network is performed on an AWS SageMaker cluster of 8 machines, each with 8 x NVIDIA A100 GPUs in $float16$ precision. Processing also includes standardized SMILES tokenization using the Tokenizers library [25]. The full search job comprises the prediction steps and top-k selection steps running on multiple machines in parallel, followed by a top-k aggregation step performed using a PostgreSQL database on a single machine (see fig. 2 for the full architecture overview). Multi-node parallel processing is an essential component of the design because reading molecular embeddings from SSD disks is a bottleneck. However, we opt out of using RAM for this task, as it would be much more expensive when keeping almost 5 TB of embeddings for Enamine REAL Space 40B in RAM. The search infrastructure is set up as follows: 27 nodes with 2-core CPUs and 4 GB RAM connected to 270 SSD disks of 21 GB each, with 1 Gbit/s of throughput per disk. With this infrastructure, we are able to cost-efficiently host clients willing to screen multi-billion molecular libraries within seconds (see section 5.5 for speed evaluation).
@@ 5.3 Hi-benchmark results
For BIOPTIC and fingerprint search, we used 10 most active compounds from the training set as queries. To get QSAR predictions from the model, we calculated cosine similarity between embeddings (or fingerprints) of queries and test set molecules. We ranked test molecules by maximal similarity to the five queries, with the most similar at the top of the ranking.
@@ 5.4 Effect of queries selection
The results are in fig. 3. All queries are selected from the training set. The "1 most active" strategy performs much worse on ACHE and EGFR, potentially due to noise in the data. The "5 most active," "10 most active," and "10 diverse active molecules" strategies are generally good.
@@ 5.5 Retrieval performance on ultra-large molecular libraries
We found this step to be memory-bound by SSD throughput, so we measured performance for the in-RAM scenario. If all the embeddings are already loaded in RAM (which might be the case to handle a lot of search queries), we can process up to 1 billion molecules per second on a single 32-core AWS Graviton3 processor, with the speed increasing with the number of cores according to the demonstration in (see fig. 4).
@@ 5.6 Scaffold hopping from the box
To demonstrate diverse candidates selection performed with BIOPTIC, we chose ACHE, AA2AR and EGFR as targets and ran search in Enamine REAL Space 40B library. We used 10 diverse active compounds as queries and retrieved top-10k candidates (selection lists). We also took 10k random compounds from Enamine REAL Space and all known molecules tested for binding to the corresponding targets from BindingDB. We then mapped Morgan fingerprints of all the molecules into 2D coordinates with t-SNE. From each selection list, we then removed all molecules that are closer than 0.4 Tanimoto similarity to any BindingDB molecule and picked 10 diverse molecules with RDKit’s MaxMinPicker. The results are shown in 5. It’s seen from the figures that BIOPTIC 1) prioritizes molecules from a widely span corner of the Enamine chemical space, 2) picks molecules structurally different from what is already known, 3) selects candidates of different scaffolds.

### v4

# BIOPTIC B1: a target-agnostic potency-based small molecules search engine

@@ 4.1 Modeling
Our model is a SMILES-based transformer [21], which we train in two phases. The Byte-Pair Encoding (BPE) tokenization [22] is applied to SMILES strings at each phase. During the first phase, the model is pre-trained in an unsupervised fashion on a large corpus of unlabeled molecular data from PubChem and Enamine REAL Space. The training is performed via a masked language modeling following the procedure from RoBERTa [23] (see Figure 1a). This phase allows the model to learn the SMILES grammar.
In the second stage, the transformer model is augmented with a multi-pooling module that aggregates hidden layer output in three different ways: taking classification token output as is, and taking maximum and average along the sequence axis of all the remaining tokens. These aggregated vectors are passed into their corresponding linear layers for projection to a lower dimension following up with concatenation to form the final embedding. This embedding is further normalized and passed to the classification layer where the task is to predict which targets a given molecule is active to. Here, the model is trained on BindingDB. As the majority of labels are not available for the BindingDB drug-target pairs, we mask them out to not introduce additional noise into the training (see Figure 1b).
@@ 4.2 Optimization
The chosen model is RoBERTa, with a vocabulary size of 500, 6 hidden layers with a hidden size of 384, 8 attention heads, and an intermediate size of 1024, while the rest of the parameters are as described in [23]. The model has a total of 8.7 million parameters. Initially, it is pre-trained in an unsupervised fashion on a large corpus of data comprising 115 million unique molecules from PubChem and a random set of 48 million molecules from the Enamine REAL Space library, totaling over 160 million molecules. The training is performed using masked language modeling following the procedure from RoBERTa [23] (Figure 1).
@@ 4.3 Ultra-fast molecules retrieval system
For the Enamine REAL Space of 40 billion molecules, the inference of the neural network is performed on an AWS SageMaker cluster of 8 machines, each with 8 x NVIDIA A100 GPUs in $float16$ precision. Processing also includes standardized SMILES tokenization using the Tokenizers library [25]. The full search job comprises the prediction steps and top-k selection steps running on multiple machines in parallel, followed by a top-k aggregation step performed using a PostgreSQL database on a single machine (see Figure 2 for the full architecture overview). Multi-node parallel processing is an essential component of the design because reading molecular embeddings from SSD disks is a bottleneck. However, we opt out of using RAM for this task, as it would be much more expensive when keeping almost 5 TB of embeddings for Enamine REAL Space 40B in RAM. The search infrastructure is set up as follows: 27 nodes with 2-core CPUs and 4 GB RAM connected to 270 SSD disks of 21 GB each, with 1 Gbit/s of throughput per disk. With this infrastructure, we are able to cost-efficiently host clients willing to screen multi-billion molecular libraries within seconds (see section 5.5 for speed evaluation).
@@ 5.3 Hi-benchmark results
For BIOPTIC and fingerprint search, we used 10 most active compounds from the training set as queries based on the analysis performed in section 5.4. To get QSAR predictions from the model, we calculated cosine similarity between embeddings (or fingerprints) of queries and test set molecules. We ranked test molecules by maximal similarity to the queries, with the most similar at the top of the ranking.
@@ 5.4 Effect of queries selection
The results are in Figure 3. All queries are selected from the training set. The "1 most active" strategy performs much worse on ACHE and EGFR, potentially due to noise in the data. The "5 most active," "10 most active," and "10 diverse active molecules" strategies are generally good.
@@ 5.5 Retrieval performance on ultra-large molecular libraries
We found this step to be memory-bound by SSD throughput, so we measured performance for the in-RAM scenario. If all the embeddings are already loaded in RAM (which might be the case to handle a lot of search queries), we can process up to 1 billion molecules per second on a single 32-core AWS Graviton3 processor, with the speed increasing with the number of cores according to the demonstration in (see Figure 4).
@@ 5.6 Scaffold hopping from the box
To demonstrate diverse candidates selection performed with BIOPTIC, we chose ACHE, AA2AR and EGFR as targets and ran search in Enamine REAL Space 40B library. We used 10 diverse active compounds as queries and retrieved top-10k candidates (selection lists). We also took 10k random compounds from Enamine REAL Space and all known molecules tested for binding to the corresponding targets from BindingDB. We then mapped Morgan fingerprints (ECFP4) of all the molecules into 2D coordinates with t-SNE. From each selection list, we then removed all molecules that are closer than 0.4 Tanimoto similarity to any BindingDB molecule and picked 10 diverse molecules with RDKit’s MaxMinPicker. The results are shown in 5. It’s seen from the figures that BIOPTIC 1) prioritizes molecules from a widely span corner of the Enamine chemical space, 2) picks molecules structurally different from what is already known, 3) selects candidates of different scaffolds.